# Rapid and Power-Aware Learned Optimization for Modular Receive Beamforming

### v1

@@ I Introduction
Massive MIMO technology is at the core of future wireless communications, with extremely large implementations expected to enable meeting the constantly growing demands in throughput and coverage [2, 3]. However, implementing communications with hundreds or thousands of antenna elements gives rise to several challenges, including issues related to cost, power consumption, deployment, and scalability [4]. To overcome these challenges, a leading approach to massive MIMO adopts modular designs. In such cases, an extremely large MIMO array is implemented using many distinct modules, each equipped with a (smaller) antenna array, and possibly possesses local beamforming capabilities. These modular architectures can take different forms, e.g., distributed cell-free massive MIMO [5], centralized interconnections of multiple panels [6], or radio stripes [7].
@@
The main task of hybrid beamforming can be viewed as the conversion of channel state information (CSI) into a beampattern achievable through a divisible architecture [18]. Various methods have been proposed for tackling this task (see, e.g., overview in [19]). As hybrid beamforming can be represented as a constrained optimization problem, a conventional approach achieves this via iterative optimizers, see, e.g., [20, 21, 22, 23]. However, these approaches typically involve lengthy optimization with numerous iterations, which can be challenging to carry out efficiently within the extremely short time frames required for determining beamforming (e.g., on the order of $0.1$ milliseconds [19]).
To facilitate rapid beamforming setting, recent works proposed using deep learning tools, taking advantage of their ability to learn to tackle challenging optimization problems with fixed and limited latency [24, 25]. This can be achieved by training conventional deep learning architectures, such as convolutional neural networks, to map MIMO CSI into analog and digital precoders [26, 27], or by combining optimization with deep learning for hybrid beamformers [28, 29, 30, 31, 32] as a form of model-based deep learning [33]. However, while hybrid beamforming can be viewed as encompassing modular beamforming, its associated algorithms are not tailored for exploiting modular structures, and do not account for the limited power considerations imposed in such forms of distributed MIMO systems. Moreover, current learning-aided designs often focus on a given configuration, and the optimizer needs to be retrained when using a different array (or adding another module), limiting scalability. Finally, existing works typically focus on downlink beamforming, with the uplink being much less studied.
@@
  • First-order optimization for modular receive beamforming: We formulate the receive modular beamforming task as a constrained optimization, and derive the corresponding first-order iterative optimizer based on PGA steps. We specifically characterize the corresponding gradient steps and projection methods, which enable (lengthy) tuning of modular receive beamformers.
  • Rapid and scalable learned optimization: To enable rapid tuning, we convert the identified PGA steps into a machine learning model operating with a fixed and low latency. This is achieved by fixing the number of iterations while using the hyperparameters of PGA with momentum as trainable machine learning parameters. We show that the interpretable operation of the resulting design allows it to be applied to different number of modules (which can be much larger than those used in training), thus facilitating scalability.
  • Power-aware learned optimization: We leverage the fact that the resulting machine learning model is comprised of interpretable PGA iterations to incorporate two forms of power-aware considerations: the ability to boost sparse modular architectures with deactivated phase shifters, and the support of low-resolution phase shifters. As the resulting constraints render PGA non-differentiable, we approximate its operation for learning purposes via surrogate projection mappings that enable its representation as a machine learning architecture.
  • Extensive experimentation: We extensively evaluate our learned optimizer for both synthetic channels as well as channels obtained from the physically compliant Quasi-Deterministic Radio channel Generator (QuaDRiGa) model [37]. Our numerical studies systematically show that learned optimization enables reliable tuning of modular beamformers with reductions of 1030×10-30$\times$10 - 30 × in iterations compared to conventional optimizers, and that the power-aware design enables efficient operation with only a minor degradation in the accuracy compared to unconstrained designs.
@@ II-B1 Panel-Wise Beamforming
Note that for all considered feasible sets, the phase shifting components remain frequency-agnostic, i.e., ${\boldsymbol{W}}_{p}$ is invariant of the frequency index $b$. Moreover, we note that approaches C2 and C3, which aim to improve the power efficiency of the panels, are complementary. Accordingly, one can restrict the analog processing at the panels side to be both sparse and quantized, i.e., have some of the phase shifters deactivated, while the active ones take discrete phase values.
@@ II-B2 Panel-CPU Connectivity
We model the connectivity to the CPU via the binary matrix ${\boldsymbol{A}}\in\{0,1\}^{L\cdot P\times T}$, with $T$ being the number of CPU inputs. This matrix is also frequency invariant, and its entry $[{\boldsymbol{A}}]_{l\cdot p,t}$ denotes whether the $l$th output of the $p$th panel is connected to the $t$th CPU input. The CPU input ${\boldsymbol{z}}[b]$ is related to the channel output ${\boldsymbol{y}}[b]$ via
@@
where ${\boldsymbol{W}}={\rm blkdiag}\{{\boldsymbol{W}}_{1},\ldots,{\boldsymbol{W}}_{P}\}$. This constraint comes from the decentralized structure of our communication system, which takes place within each panel. Note that when ${\boldsymbol{A}}$ is the identity matrix and $L=1$, this formulation specializes the common form of partially-connected hybrid beamformers [11].
@@ II-C Problem Formulation
We aim to design the modular receive beamformer detailed in (5) to maximize the achievable rate. To formulate this, we define the equivalent channel at the $b$th frequency bin as ${\boldsymbol{G}}\triangleq{\boldsymbol{W}}{\boldsymbol{A}}\in\mathbb{C}^{M\times T}$. For the uplink setup detailed in Subsection II-A, the average achievable sum-rate is given by (see, e.g., [39, Thm. 1])
@@
Using the rate expression in (6), we can formulate the setup of modular receive beamforming as a constrained optimization problem. Since the matrix ${\boldsymbol{A}}$ represents the hardware connectivity between the panels, it is assumed to be constant. Accordingly, for a given channel realization $\{{\boldsymbol{H}}[b]\}_{b=1}^{B}\}$, the modular beamforming task is formulated as the recovery of
@@ III-B Unfolded PGA with Momentum
Algorithm 1 allows tackling (7) while meeting the requirement to operate with $J$ iterations. However, iterative optimizers such as PGA are designed to tackle convex optimization problems and to operate with asymptotically large number of iterations. There, convergence can be guaranteed with fixed hyperparameters $\{\mu,\beta\}$ (i.e., independent of the iteration index $j$), which can be tuned manually. In Algorithm 1, we apply this convex optimizer to a non-convex problem (7) while operating with a fixed and small number of iterations $J$. In such cases, the setting of the hyperparameters $\{\mu_{j},\beta_{j}\}$ greatly affects performance [40, Ch. 9]. This motivates leveraging the available data in (8) to tune these hyperparameters, by converting Algorithm 1 into a discriminative machine learning model [41], and proposing a dedicated learning mechanism. We next detail the resulting machine learning architecture, its proposed training procedure, and how it can be scaled to large modular settings.
@@ III-B2 Training
Converting Algorithm 1 into a machine learning model facilitates tuning ${\boldsymbol{\theta}}$ in (11), which are the hyperparameters of Algorithm 1. This is achieved using the dataset $\mathcal{D}$ in (8), containing past channel realizations taken from the same communication system. Letting ${\boldsymbol{W}}_{J}({\boldsymbol{\theta}})$ be the modular beamforming matrix computed by Algorithm 1 with parameters ${\boldsymbol{\theta}}$, we seek the hyperparameters
${\boldsymbol{\theta}}^{\star}=\mathop{\arg\max}\limits_{{\boldsymbol{\theta}}}\sum_{\{{\boldsymbol{H}}[b]\}\in\mathcal{D}}R({\boldsymbol{W}}_{J}({\boldsymbol{\theta}}){\boldsymbol{A}};\{{\boldsymbol{H}}[b]\}).$ (12)
@@
$\mathcal{L}_{\mathcal{D}}({\boldsymbol{\theta}})=-\sum_{\{{\boldsymbol{H}}[b]\}\in\mathcal{D}}\sum_{j=1}^{J}\frac{\log(1+j)}{|\mathcal{D}|}R({\boldsymbol{W}}_{j}({\boldsymbol{\theta}}){\boldsymbol{A}};\{{\boldsymbol{H}}[b]\}).$
@@ 4             Apply model with parameters ${\boldsymbol{\theta}}$ to $\mathcal{D}_{q}$;
III-B3 Transferability Across Configurations
@@ IV-A Sparsified Modular Beamformer
Due to the inherent challenges associated with gradient-based optimization under $\ell_{0}$ constraints, we adopt the common approach of relaxing it into an $\ell_{1}$ constraint (see, e.g., [46, Ch. 1]). The relaxed optimization problem is
@@ IV-C Power Aware PGA
Algortihm 3 is designed to tackle the maximization problems outlined in (16) and (19) by proper settings of the projection operator and the regularization coefficient. For instance, by setting the projection to be given in (20) and setting $\lambda=0$, Algorithm 3 tackles problem (19). Similarly, by using $\lambda>0$ and setting the projection to be (18), it tackles problem (16).
@@ IV-D2 Projection
    In (22) $\sigma$ denoting the sigmoid function, $s_{m}$ denotes a sharpness hyperparameter, and $\zeta_{m}$ stands for the shift, imitating the threshold $\zeta$ from (18).
@@ IV-D3 Loss Function
${\boldsymbol{\theta}}^{\star}=\mathop{\arg\max}\limits_{{\boldsymbol{\theta}}}\sum_{\{{\boldsymbol{H}}[b]\}\in\mathcal{D}}R({\boldsymbol{W}}_{J}({\boldsymbol{\theta}}){\boldsymbol{A}};\{{\boldsymbol{H}}[b]\})$ (24)
@@
$\mathcal{L}_{\mathcal{D}}({\boldsymbol{\theta}})=-\frac{1}{|\mathcal{D}|}\sum_{\{{\boldsymbol{H}}[b]\}\in\mathcal{D}}$ $\sum_{j=1}^{J}\log(1+j)R({\boldsymbol{W}}_{j}({\boldsymbol{\theta}}){\boldsymbol{A}};\{{\boldsymbol{H}}[b]\})$
$-\gamma(\|{\boldsymbol{W}}_{J}({\boldsymbol{\theta}})\|_{1}-S_{\max})^{2},$ (25)

### v2

@@ I Introduction
Massive MIMO technology is at the core of future wireless communications, with extremely large implementations expected to enable meeting the constantly growing demands in throughput and coverage [2, 3]. However, implementing communications with hundreds or thousands of antenna elements gives rise to several challenges, including issues related to cost, power consumption, deployment, and scalability [4]. To overcome these challenges, a leading approach to massive MIMO adopts modular designs. In such cases, a MIMO array is implemented using many distinct modules, each equipped with a (smaller) antenna array, and possibly possesses local beamforming capabilities. These architectures encompass various forms of emerging MIMO technologies, including extremely large arrays [5, 6], modular low-frequency deployments [7] distributed cell-free massive MIMO [8], centralized interconnections of multiple panels [9, 10], and even radio stripes [11].
@@
The main task of hybrid beamforming can be viewed as the conversion of channel state information (CSI) into a beampattern achievable through a divisible architecture [22]. Various methods have been proposed for tackling this task (see, e.g., overview in [23]). As hybrid beamforming can be represented as a constrained optimization problem, a conventional approach achieves this via iterative optimizers, see, e.g., [24, 25, 26]. However, these approaches typically involve lengthy optimization with numerous iterations, which can be challenging to carry out efficiently within the extremely short time frames required for determining beamforming (e.g., on the order of $0.1$ milliseconds [23]).
To facilitate rapid beamforming setting, recent works proposed using deep learning tools, taking advantage of their ability to learn to tackle challenging optimization problems with fixed and limited latency [27, 28]. This can be achieved by training conventional deep learning architectures, such as convolutional neural networks, to map MIMO CSI into analog and digital precoders [29, 30], or by combining optimization with deep learning for hybrid beamformers [31, 32, 33, 34] and for fully digital beamformers [35] as a form of model-based deep learning [36]. However, while hybrid beamforming can be viewed as encompassed by modular beamforming, its associated algorithms are not tailored for exploiting modular structures, and do not account for the limited power considerations imposed in such forms of distributed MIMO systems. Moreover, current learning-aided designs often focus on a given configuration, and the optimizer needs to be retrained when using a different array (or adding another module), limiting scalability. Finally, existing works typically focus on downlink beamforming, with the uplink being much less studied.
@@
  • First-order optimization for modular receive beamforming: We formulate the receive modular beamforming task as a constrained optimization, and derive the corresponding first-order iterative optimizer based on PGA steps. The resulting optimization formulation differs from previous works that focused on the downlink, namely, transmit beamforming. We specifically characterize the corresponding gradient steps and projection methods, which enable (lengthy) tuning of modular receive beamformers.
  • Rapid and scalable learned optimization: To enable rapid tuning, we convert the identified PGA steps into a machine learning model operating with a fixed and low latency. This is achieved by fixing the number of iterations while using the hyperparameters of PGA with momentum as trainable machine learning parameters. We show that the interpretable operation of the resulting design allows it to scale to different numbers of modules (potentially much larger than those used during training), thus facilitating scalability, and that its training procedure can be adapted to enhance robustness to noisy channel estimates.
  • Power-aware learned optimization: We leverage the interpretable structure of the resulting machine learning model, comprised of PGA iterations, to incorporate two forms of power-aware considerations: the ability to boost sparse modular architectures with deactivated phase shifters, and the support of low-resolution phase shifters. As the resulting constraints render PGA non-differentiable, we approximate its operation for learning purposes via surrogate projection mappings that enable its representation as a machine learning architecture.
  • Extensive experimentation: We extensively evaluate our learned optimizer on both synthetic channels and channels obtained from the physically compliant Quasi-Deterministic Radio channel Generator (QuaDRiGa) model [40]. Our numerical studies systematically show that learned optimization enables reliable tuning of modular beamformers with reductions of 715×7-15$\times$7 - 15 × in iterations compared to conventional optimizers. Moreover, we show that the power-aware design enables efficient operation, typically with only a minor degradation in accuracy compared to unconstrained designs. It is also shown that under some forms of limited designs, our method achieves superior performance in just $10$ iterations, significantly outperforming other methods, which required $500$ iterations to converge to a suboptimal solution. Additionally, the proposed method exhibits improved robustness under CSI estimation errors, demonstrating a notably softer degradation in performance compared to traditional approaches as the CSI noise variance increases. These results reveal the versatility of our approach, which balances computational efficiency and accuracy, even under challenging power, hardware, or imperfect CSI conditions.
@@ II-B1 Panel-Wise Beamforming
Note that for all considered feasible sets, the phase shifting components remain frequency-agnostic, i.e., ${\boldsymbol{W}}_{p}$ is invariant of the frequency index $b$. It is also noted that the constrained form of receive beamforming in general cannot realize conventional simple receive beamformers, such as the zero-forcing of minimal mean squared-error beamformers [42, Ch. 8.3]. Moreover, we note that approaches C2 and C3, which aim to improve the power efficiency of the panels, are complementary. Accordingly, one can restrict the analog processing at the panels side to be both sparse and quantized, i.e., have some of the phase shifters deactivated, while the active ones take discrete phase values.
@@ II-B2 Panel-CPU Connectivity
We model the connectivity to the CPU via the binary matrix ${\boldsymbol{A}}\in\{0,1\}^{L\cdot P\times T}$, with $T$ being the number of CPU inputs. This matrix is also frequency invariant, and its entry $[{\boldsymbol{A}}]_{l\cdot p,t}$ denotes whether the $l$th output of the $p$th panel is connected to the $t$th CPU input. The matrix ${\boldsymbol{A}}$ represents the physical connections (e.g., wiring) between the panels and the centralized processor, and is thus assumed to fixed by the modular MIMO infrastructure. The CPU input ${\boldsymbol{z}}[b]$ is related to the channel output ${\boldsymbol{y}}[b]$ via
@@
where ${\boldsymbol{W}}={\rm blkdiag}\{{\boldsymbol{W}}_{1},\ldots,{\boldsymbol{W}}_{P}\}$. This constraint comes from the decentralized structure of our communication system, which takes place within each panel.
The above model for modular hybrid receive processing accommodates both distributed MIMO settings with separated panels [8], as well as massive MIMO architectures where the panels are collocated on the same platform [10]. The distinction between such settings lies in the resulting channel matrices $\{{\boldsymbol{H}}[b]\}$, which we do not constrain to take a specific form. Moreover, our model can be viewed as specializing different forms of hybrid MIMO systems when ${\boldsymbol{A}}$ is the identity matrix, e.g., having $L=1$ in such cases specializes the common form of partially-connected hybrid beamformers, while setting $P=1$ realizes fully-connected hybrid beamformers [15].
@@ II-C Problem Formulation
We aim to design the modular receive beamformer detailed in (5) to maximize the achievable rate. To formulate this, we define the equivalent channel at the $b$th frequency bin as ${\boldsymbol{G}}\triangleq{\boldsymbol{W}}{\boldsymbol{A}}\in\mathbb{C}^{M\times T}$. For the uplink setup detailed in Subsection II-A, the average achievable sum-rate is given by (see, e.g., [43, Thm. 1])
@@
Using the rate expression in (6), we can formulate the setup of modular receive beamforming as a constrained optimization problem. Since the matrix ${\boldsymbol{A}}$ represents the hardware connectivity between the panels, it is assumed to be constant, thus not optimized. Accordingly, for a given channel realization $\{{\boldsymbol{H}}[b]\}_{b=1}^{B}\}$, the modular beamforming task is formulated as the recovery of
@@ III-B Unfolded PGA with Momentum
Algorithm 1 allows tackling (7) while meeting the requirement to operate with $J$ iterations. However, iterative optimizers such as PGA are designed to tackle convex optimization problems and to operate with asymptotically large number of iterations. There, convergence can be guaranteed with fixed hyperparameters $\{\mu,\beta\}$ (i.e., independent of the iteration index $j$), which can be tuned manually. In Algorithm 1, we apply this convex optimizer to a non-convex problem (7) while operating with a fixed and small number of iterations $J$. In such cases, the setting of the hyperparameters $\{\mu_{j},\beta_{j}\}$ greatly affects performance [44, Ch. 9]. This motivates leveraging the available data in (8) to tune these hyperparameters, by converting Algorithm 1 into a discriminative machine learning model [45], and proposing a dedicated learning mechanism. We next detail the resulting machine learning architecture, its proposed training procedure, how its trainability can leveraged to increase robustness to CSI errors, and how it can scaled to large modular settings.
@@ III-B2 Training
Converting Algorithm 1 into a machine learning model facilitates tuning ${\boldsymbol{\theta}}$ in (11), which are the hyperparameters of Algorithm 1. This is achieved using the dataset $\mathcal{D}$ in (8), containing past channel realizations taken from the same communication system. Letting ${\boldsymbol{W}}_{J}({\boldsymbol{\theta}};\{{\boldsymbol{H}}[b]\})$ be the modular beamforming matrix computed by Algorithm 1 with parameters ${\boldsymbol{\theta}}$ when applied to channel $\{{\boldsymbol{H}}[b]\}$, we seek the hyperparameters
$\!{\boldsymbol{\theta}}^{\star}=\!\mathop{\arg\max}\limits_{{\boldsymbol{\theta}}}\sum_{\{{\boldsymbol{H}}[b]\}\in\mathcal{D}}\!R({\boldsymbol{W}}_{J}({\boldsymbol{\theta}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0};\{{\boldsymbol{H}}[b]\}}){\boldsymbol{A}};\{{\boldsymbol{H}}[b]\}).$ (12)
@@
$\mathcal{L}_{\mathcal{D}}({\boldsymbol{\theta}})=-$$\sum_{\{{\boldsymbol{H}}[b]\}\in\mathcal{D}}\sum_{j=1}^{J}\frac{\log(1+j)}{|\mathcal{D}|}$
×$\qquad\times R({\boldsymbol{W}}_{j}({\boldsymbol{\theta}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0};\{{\boldsymbol{H}}[b]\}}){\boldsymbol{A}};\{{\boldsymbol{H}}[b]\}).$ (13)
@@ 4             Apply model with parameters ${\boldsymbol{\theta}}$ to $\mathcal{D}_{q}$;
III-B3 CSI Estimation Error Robustness
The iterative optimizer presented in this work assumes perfect CSI. However, in practice one may only have access to a noisy estimate of the CSI. The fact that our unfolded PGA is trained from data allows it to also to learn hyperparameters that boost robustness to such noisy CSI, without altering the operation of the optimization algorithm.
The presence of CSI errors leads to having the optimizer applied to perturbed channel realizations, represented as ${{\boldsymbol{H}}[b]+{\boldsymbol{E}}[b]}$, rather than the true channel realizations ${{\boldsymbol{H}}[b]}$. Here, ${\boldsymbol{E}}[b]$ denotes the CSI estimation error. The trainability of unfolded PGA allows it to be trained with noisy CSI to yield beampatterns that are suitable for the true CSI. To that aim, we model the CSI error as an i.i.d. Gaussian noise matrix with zero mean and variance $\sigma_{E}^{2}$, which represents the expected level of estimation error. During training, where one has access to the channel realizations as in 8, instead of applying the trainable optimizer to the true channel realizations, we apply it to a noisy version of the channels. To compute the loss during training, the noiseless channels ${{\boldsymbol{H}}[b]}$ are used to calculate the rate. This approach ensures that the optimizer learns hyperparameters that yield beamformers tailored to the true underlying channels, rather than to their noisy estimates.
Mathematically, the loss function in (III-B2) is updated as follows:
$\mathcal{L}_{\mathcal{D}}^{\sigma_{E}^{2}}({\boldsymbol{\theta}})=$$-\sum_{\{{\boldsymbol{H}}_{i}[b]\}\in\mathcal{D}}\sum_{j=1}^{J}\frac{\log(1+j)}{|\mathcal{D}|}$
×$\times R({\boldsymbol{W}}_{j}({\boldsymbol{\theta}};\{{\boldsymbol{H}}_{i}[b]+{\boldsymbol{E}}_{i}[b]\}){\boldsymbol{A}};\{{\boldsymbol{H}}_{i}[b]\}),$ (14)
where ${\boldsymbol{E}}_{i}[b]\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(0,\sigma_{E}^{2})$ is sampled anew for each forward pass during training. When $\sigma_{E}^{2}=0$, the training procedure reduces to the noiseless CSI case detailed in Subsection III-B2. The resulting noisy CSI aware training allows achieving improved robustness via hyperparameter learning, as empirically demonstrated in Subsection V-B3.
III-B4 Transferability Across Configurations
@@ III-C Discussion
Our formulation of the modular hybrid beamforming setting considers the connectivity matrix ${\boldsymbol{A}}$ to be fixed, representing a given physical infrastructure or wiring. Still, there are scenarios where it is of interest to also optimize ${\boldsymbol{A}}$. These including the possibility of optimizing it for each channel alongside ${\boldsymbol{W}}$, e.g., when the hardware includes rapidly adaptable switches, or even for having a single ${\boldsymbol{A}}$ optimized for multiple coherence duration, thus effectively guiding the system design. While both options can be studied based on our proposed learned optimization methodology, we leave these extensions of our work for future study.
@@ IV-A Sparsified Modular Beamformer
Due to the inherent challenges associated with gradient-based optimization under $\ell_{0}$ constraints, we adopt the common approach of relaxing it into an $\ell_{1}$ constraint (see, e.g., [50, Ch. 1]). The relaxed optimization problem is
@@ IV-C Power Aware PGA
Algorithm  3 is designed to tackle the maximization problems outlined in (18) and (21) by proper settings of the projection operator and the regularization coefficient. For instance, by setting the projection to be given in (22) and setting $\lambda=0$, Algorithm 3 tackles problem (21). Similarly, by using $\lambda>0$ and setting the projection to be (20), it tackles problem (18).
@@ IV-D2 Projection
    In (24), $\sigma$ denotes the sigmoid function, $s_{m}$ denotes a sharpness hyperparameter, and $\zeta_{m}$ stands for the shift, imitating the threshold $\zeta$ from (20).
@@ IV-D3 Loss Function
${\boldsymbol{\theta}}^{\star}=\mathop{\arg\max}\limits_{{\boldsymbol{\theta}}}\sum_{\{{\boldsymbol{H}}[b]\}\in\mathcal{D}}R({\boldsymbol{W}}_{J}({\boldsymbol{\theta}};{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\{{\boldsymbol{H}}[b]\}}){\boldsymbol{A}};\{{\boldsymbol{H}}[b]\})$ (26)
@@
$\mathcal{L}_{\mathcal{D}}({\boldsymbol{\theta}})=$
$-\frac{1}{|\mathcal{D}|}\sum_{\{{\boldsymbol{H}}[b]\}\in\mathcal{D}}\sum_{j=1}^{J}\log(1+j)R({\boldsymbol{W}}_{j}({\boldsymbol{\theta}};{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\{{\boldsymbol{H}}[b]\}}){\boldsymbol{A}};\{{\boldsymbol{H}}[b]\})$
$-\gamma(\|{\boldsymbol{W}}_{J}({\boldsymbol{\theta}};{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\{{\boldsymbol{H}}[b]\}})\|_{1}-S_{\max})^{2},$ (27)